# Systematic search for cool hypervelocity stars in multi-epoch WISE data

### v1

@@ Abstract
Context. Sixty years after the discovery of brown dwarfs, the search for these objects continues, particularly in the vicinity of the Sun. Objects near the Sun are characterized by large proper motions, making them seen as fast-moving objects. While the Gaia DR3 catalogue is a comprehensive source of proper motions, it lacks the depth needed for discovering fainter objects. Modern multi-epoch surveys, with their greater depth, offer a new opportunity for systematic search for ultra-cool dwarfs.
@@
Methods. We used recently released unTimely catalogue of epochal detections in unWISE coadds to search for objects with high proper motions using simple motion detection algorithm. This method was used to identify objects with proper motions exceeding approximately 0.6 arcseconds per year. The identified objects were then cross-referenced with data from other large-scale sky surveys to further analyze their characteristics.
Results. The search yielded 3245 moving objects with significant proper motions, 32 of which had not been previously published. Among these, at least 15 were identified as reliable new brown dwarf candidates, with estimated distances closer than 50 parsecs and spectral types later than T0.
@@
According to recent studies (Mužić et al., 2017), the number of brown dwarfs in the Galaxy ranges from 25 to 100 billion objects (with the total number of objects ranging from 100 to 500 billion). Homogeneous and complete samples of brown dwarfs are needed for various kinds of studies: kinematic studies (Smith et al., 2014), studies of binary stars with brown dwarfs (Lodieu et al., 2014), and studies of the parameters of the Galaxy.
@@
Such selection, however, was rarely perfect, even with a combination of a number of surveys. Photometric samples are often contaminated with extraneous objects, requiring spectroscopic confirmation to establish reliable samples. The only search for brown dwarfs based on spectra was conducted by (Schmidt et al., 2010). The work utilised SDSS DR7 spectroscopic database and provided 210 newly discovered brown dwarfs, showing that the existing color selection method are biased towards redder brown dwarfs.
@@
As brown dwarfs are very faint and therefore only can be found close to the Sun, their proper motions should be large, compared to the majority of the field stars. Combined with typically suboptimal angular resolution of infrared instruments, it significantly complicates the association of measurements between different datasets acquired at different epochs, leading to predominantly slower moving objects being detected using color selection methods.
@@
To overcome this bias, uniform data sets with dense multi-epoch coverage are needed. The multi-epoch data from 2MASS survey (Kirkpatrick et al., 2010) and from AllWISE Motion Survey (Kirkpatrick et al., 2014) with a spectroscopic follow-up have revealed previously unknown LT dwarfs and subdwarfs. The large-scale search for rapidly moving brown dwarfs is nowadays enabled by the CatWISE (Marocco et al., 2021) catalogue that reports both infrared color and proper motions for nearly two billions of objects over the whole sky. However, as the object detection in CatWISE is performed on the co-added images combining all unWISE epochs, it is not optimal for finding rapidly moving stars, and prone to numerous spurious detections due to aggressive deblending applied to the data. Thus, the projects like Backyard Worlds have to utilize citizen science approach for visual inspection of many candidate sources in order to find actual moving objects (Kuchner et al., 2017; Rothermich et al., 2024; Marocco et al., 2024).
@@
On the other hand, recent release of unTimely Meisner et al. (2023) catalogue of epochal detections in individual unWISE coadds opens the possibility of direct search for most rapidly moving objects. Here we describe such a search that we performed on W2 band data in order to specifically find new candidates for nearby cool dwarfs. In Section 2 we describe the algorithm and data we used, as well as the moving objects we detected. Section 3 contains the discussion of the sensitivity of our search, presents the list of high proper motion objects we detected that lack both Gaia detections and Simbad records, and reviews their properties and possible classification as late-type cool dwarf candidates. Finally, Section 4 concludes the work.
@@ 2 Data and analysis
The Wide-field Infrared Survey Explorer, or WISE (Wright et al., 2010), is a space telescope that performed all-sky survey initially in four infrared bands (W1 at 3.4m, W2 at 4.6m, W3 at 12m and W4 at 22m), and since the coolant depletion in late 2010 – in two shorter-wavelengths bands only. It repeatedly scanned the sky in great circles near a solar elongation of 90, typically observing a given region of the sky over a period of 1 d every 6 months, with denser coverage closer to the ecliptic poles. unWISE coadds combine the exposures from the same sky regions into a series of six-monthly visits, with an average of typical 12 exposures per band per visit (Meisner et al., 2019, 2018). These coadds are both sharper and deeper than original data used for producing initial AllWISE catalogue (Cutri et al., 2021), and are the basis of unWISE (Schlafly et al., 2019) and CatWISE2020 (Marocco et al., 2021) catalogues greatly improving over it. However, these catalogues still did not provide the detections or measurements for the objects at individual epochs.
@@
In order to systematically search for cool rapidly moving objects in WISE data we used recently released unTimely catalogue (Meisner et al., 2023) which is the result of uniform analysis of individual epochal unWISE coadds (Meisner et al., 2019, 2018) using crowdsource crowded-field photometry package (Schlafly et al., 2018; Schlafly, 2021). It is currently distributed as a set of FITS tables each corresponding to the objects detected in one of two filters (W1 or W2) in individual epochal coadds for one of 18240 sky tiles. Typically every tile is covered with 15 to 17 unWISE coadds spanning approximately 11 years between Jan 2010 and Dec 2020 in every filter, but some of them (284 tiles, located close to ecliptic poles) have significantly more, up to two hundred, epochs spanning the same time interval (Meisner et al., 2018).
@@
Epochal catalogues are independent, and not cross-matched together, both between different epoch and between different filters in a single epoch. Therefore, as we are primarily interested in detecting cool objects, we processed only the catalogues in W2 (4.6m) band. Moreover, to ensure the consistency of depths of catalogues from individual epochs, we excluded from the analysis the epochs with total number of contributing exposures (N_EXP field in the unTimely metadata table) less than a half of the median value for that tile. We applied basic quality cuts to the objects from epochal catalogues by selecting only the detections from primary parts of unWISE coadds (primary==1) having flux measured with signal to noise ratio of at least 3. Additionally, we require the detections to have sufficiently large fraction of flux inside the PSF that comes from this object (fracflux¿0.5) to ensure that the detections are not subject to significant crowding, and thus have reliable astrometry and (less relevant for us) photometry. After initial experimenting we decided not to perform any filtering on “quality factor” (qf) field or flags derived either from unWISE coadds themselves (flags_unwise) or from crowdsource processing (flags_info) as they prevent most brighter objects from being detected, while not reducing significantly the amount of typical artefacts among the candidates.
@@ 2.1 Motion detection algorithm
After applying various quality cuts as described above we analyzed the sets of all objects from epochal catalogues for every tile in order to find candidate moving objects. To do so, we converted their sky coordinates to Cartesian ones on the unit sphere, and searched for linear structures in the resulting four-dimensional space (three spatial coordinates plus time). We used a simple “opportunistic” algorithm that consider all pairs of points from different epochs that satisfy the criteria of being separated by at least 6′′ (slightly more than 2 pixels in unWISE coadds, and slightly less than full-width at half maximum (FWHM) of its point-spread function (PSF)) but no more than 90′′ (for reducing the computational complexity), and not belonging to clusters containing more objects than half of number of epochs, within 1′′. The latter criterion helps filtering out stationary objects, while the former one ensures significant motion of the candidates, with both placing a lower limit on the detectable proper motion of about 0.6′′/year111On the other hand, there is no practical upper limit on detectable proper motion as long as the track remains approximately linear.. We also excluded the pairs that contain the points already claimed by previously selected candidates – it helps uniquely identifying the candidate tracks by their first point, but, potentially, complicates the detection in too crowded fields. Then, for every pair, we formed a candidate track, computed its expected positions for all the epochs available for that tile, and then selected the objects from corresponding epochs (this time without applying any additional criteria on top of initial quality cuts) that are closer than 1′′ to these positions222While this radius of 1′′ is significantly larger than typical positional uncertainties claimed in the catalogue we decided to adopt it and deal with spurious associations later on the trajectory fitting stage.. We rejected the candidate tracks that do not have objects in more than 60% of the epochs, corresponding to at least 10 epochs in typical 16-epochal tile.
@@
The candidates selected this way are prone to both genuine image-level artefacts, especially due to long spikes around bright stars, and spurious tracks formed from individual detections of stationary objects located approximately along a line. We could not find a way to reliably filter out either kind of artefacts using the information available in the catalogue files or positional residuals around the linear fit, and therefore visually checked all candidates with a simple dashboard based on the WISEView(Meisner et al., 2018; Caselden et al., 2018) cutout API developed in the frame of Backyard Worlds project(Kuchner et al., 2017). The dashboard displays, for every candidate, the cutouts from the first and last epochs at the object position, and the corresponding differences between these images and the median image of all epochs, and allows easy selection of real moving objects by the characteristic dipoles in the latter images, as shown in Figure 1. In the unclear cases, we also directly checked the epochal animations using the original WISEView web tool333Available at http://byw.tools/wiseview..
This way, we selected 3245 candidates that passed visual inspection and thus reliably correspond to actual moving objects. Figure 2 shows the distribution of number of epochs where the object is detected, for both final candidates and artefacts. The amount of the latters significantly grows towards lower number of epochs, thus justifying limits on the selection parameters we used for the motion detection above, meaning that further releasing them is unpractical for this specific “opportunistic” algorithm.
2.2 Proper motions
For the final candidates that passed visual inspection we computed the proper motion by performing robust linear regression of their standard coordinates versus time using RANSAC(Fischler & Bolles, 1981) algorithm as implemented in scikit-learn(Pedregosa et al., 2011) Python package in order to mitigate the influence of spurious associations of stationary objects with the track due to too permissive matching radius we used. Knowing proper motion, we then performed a cross-match with Gaia DR3 catalogue (Gaia Collaboration et al., 2023) by propagating object positions to its J2016.0 epoch, using 2′′ matching radius, and only considering Gaia objects with significant proper motions to avoid confusions with faint stationary background stars. This way we found the associations for 2689 (82%) of all candidates. Left panel of Figure 3 shows that the agreement of proper motions is sufficiently good and unbiased for its whole range, while Figure 4 shows that the fraction of unassociated objects does not depend on the proper motion value, while slightly increasing for fainter fluxes444We should also note that some of unassociated objects, especially brighter ones, do actually have matching counterparts in either Gaia DR2 or Gaia eDR3 catalogues..

### v2

# New ultracool dwarf candidates from multi-epoch WISE data

@@ Abstract
Context. Thirty years after the discovery of brown dwarfs, the search for these objects continues, particularly in the vicinity of the Sun. Objects near the Sun are characterized by large proper motions, making them seen as fast-moving objects. While the Gaia DR3 catalogue is a comprehensive source of proper motions, it lacks the depth needed for discovering fainter objects. Modern multi-epoch surveys, with their greater depth, offer a new opportunity for systematic search for ultracool dwarfs.
@@
Methods. We used recently released unTimely catalogue of epochal detections in unWISE coadds to search for objects with high proper motions using simple motion detection algorithm, combined with machine learning based artefact rejection routine. This method was used to identify objects with proper motions exceeding approximately 0.3 arcseconds per year. The identified objects were then cross-referenced with data from other large-scale sky surveys to further analyze their characteristics.
Results. The search yielded 21,885 moving objects with significant proper motions, 258 of which had not been previously published. All except 6 of them are compatible with being ultracool dwarfs. Among these, at least 33 were identified as most promising new T dwarf candidates, with estimated distances closer than about 40 parsecs, and effective temperatures less than 1300 K.
@@
According to the studies (Mužić et al., 2017), brown dwarfs are estimated to constitute up to 25% of the stellar and substellar objects in the Galaxy. Recent studies (Kirkpatrick et al., 2024) offer more conservative estimates, suggesting a star-to-brown dwarf ratio of 4:1. Searching for brown dwarfs is crucial for understanding the full spectrum of stellar and substellar populations in the Galaxy. Brown dwarfs are needed for advancing our understanding of the initial mass function (IMF), a key astrophysical tool for exploring star formation, the evolution of galaxies, and the distribution of stellar masses (Chabrier, 2003; Kroupa et al., 2013; Raghu et al., 2024; 2024ApJS..271...55K; Kroupa et al., 2024). Finding new brown dwarfs enhances our understanding of Galactic kinematics and expands our understanding of stellar populations (Kirkpatrick et al., 2010; Smith et al., 2014; Luhman, 2024). Additionally, investigating the atmospheres of brown dwarfs is vital for unraveling the complex processes that influence their temperature, composition, and cloud formation (Khandrika et al., 2013; Burningham et al., 2017), ultimately providing a deeper understanding of similar mechanisms in the atmospheres of gas giant exoplanets (Marley & Robinson, 2015; Charnay et al., 2018; Tan & Showman, 2019; Tremblin et al., 2019).
@@
Such selection, however, has rarely been perfect, even with a combination of various surveys. Photometric samples are often contaminated with extraneous objects, necessitating spectroscopic confirmation to establish reliable datasets. While the search conducted by Schmidt et al. (2010) utilized the SDSS DR7 spectroscopic database and resulted in the discovery of 210 new brown dwarfs, revealing that existing color selection methods are biased towards redder brown dwarfs, it is important to acknowledge that this was not the only spectroscopic search for brown dwarfs. Subsequent studies, such as Wang et al. (2022), who explored UCD spectra in LAMOST data, and Hsu et al. (2024), who performed a comprehensive analysis of UCD candidates in APOGEE DR17, have also contributed significantly to the field. Additionally, Langeveld et al. (2024) utilized JWST/NIRISS to identify brown dwarfs in NGC 1333.
@@
Multiple brown dwarfs have been recently discovered at kiloparsec distances, see the results from JADES/CEERS surveys Hainline et al. (2024), and UNCOVER Burgasser et al. (2024a) survey. Furthermore, brown dwarfs have been recently discovered in closest globular clusters Gerasimov et al. (2024). However, most brown dwarfs are discovered in the nearest solar neighbourhood, so their proper motions should be large, compared to the majority of the field stars. Combined with typically suboptimal angular resolution of infrared instruments, it significantly complicates the association of measurements between different datasets acquired at different epochs, leading to predominantly slower moving objects being detected using color selection methods.
@@
To overcome this bias, uniform data sets with dense multi-epoch coverage are needed. The multi-epoch data from 2MASS survey (Kirkpatrick et al., 2010) and from AllWISE Motion Survey (Kirkpatrick et al., 2014) with a spectroscopic follow-up have revealed previously unknown LT dwarfs and subdwarfs. The large-scale search for rapidly moving brown dwarfs is nowadays enabled by the CatWISE2020 (Marocco et al., 2021) catalogue that reports both infrared color and proper motions for nearly two billions of objects over the whole sky. However, as the object detection in CatWISE2020 is performed on the co-added images combining all unWISE epochs, it is not optimal for finding rapidly moving stars, and prone to numerous spurious detections due to aggressive deblending applied to the data. Thus, the projects like Backyard Worlds have to utilize citizen science approach for visual inspection of many candidate sources in order to find actual moving objects (Kuchner et al., 2017; Rothermich et al., 2024; Marocco et al., 2024), as well as advanced machine learning based selection of brown dwarf candidates (Humphreys et al., 2020; Kota et al., 2022).
@@
On the other hand, recent release of unTimely (Meisner et al., 2023) catalogue of epochal detections in individual unWISE coadds opens the possibility of direct search for most rapidly moving objects. Here we describe such a search that we performed on band data in order to specifically find new candidates for nearby ultracool dwarfs, i.e. late M stars and brown dwarfs. In Section 2 we describe the algorithm and data we used, as well as the moving objects we detected. Section 3 contains the comparison of our method performance to other brown dwarf search efforts, reviews the properties and possible classification of new ultracool dwarf candidates we found. Finally, Section 4 concludes the work.
@@ 2 Data and analysis
The Wide-field Infrared Survey Explorer, or WISE (Wright et al., 2010), is a space telescope that performed all-sky survey initially in four infrared bands ( at 3.4m, at 4.6m, at 12m and at 22m), and since the coolant depletion in late 2010 – in two shorter-wavelengths bands only. It repeatedly scanned the sky in great circles near a solar elongation of 90, typically observing a given region of the sky over a period of 1 d every 6 months, with denser coverage closer to the ecliptic poles. unWISE coadds combine the exposures from the same sky regions into a series of six-monthly visits, with an average of typical 12 exposures per band per visit (Meisner et al., 2019, 2018). These coadds are both sharper and deeper than original data used for producing initial AllWISE catalogue (Cutri et al., 2021), and are the basis of unWISE (Schlafly et al., 2019) and CatWISE2020 (Marocco et al., 2021) catalogues greatly improving over it. However, these catalogues still did not provide the detections or measurements for the objects at individual epochs.
@@
In order to systematically search for cool rapidly moving objects in WISE data we used recently released unTimely catalogue (Meisner et al., 2023) which is the result of uniform analysis of individual epochal unWISE coadds (Meisner et al., 2019, 2018) using crowdsource crowded-field photometry package (Schlafly et al., 2018; Schlafly, 2021). It is currently distributed as a set of FITS tables each corresponding to the objects detected in one of two filters ( or ) in individual epochal coadds for one of 18240 sky tiles. Typically every tile is covered with 15 to 17 unWISE coadds spanning approximately 11 years between Jan 2010 and Dec 2020 in every filter, but some of them (284 tiles, located close to ecliptic poles) have significantly more, up to two hundred, epochs spanning the same time interval (Meisner et al., 2018).
@@
Epochal catalogues are independent, and not cross-matched together, both between different epoch and between different filters in a single epoch. Therefore, as we are primarily interested in detecting cool objects, we processed only the catalogues in (4.6m) band. Moreover, to ensure the consistency of depths of catalogues from individual epochs, we excluded from the analysis the epochs with total number of contributing exposures (N_EXP field in the unTimely metadata table) less than a half of the median value for that tile. We applied basic quality cuts to the objects from epochal catalogues by selecting only the detections from primary parts of unWISE coadds (primary==1) having flux measured with signal to noise ratio of at least 5. Additionally, we require the detections to have sufficiently large fraction of flux inside the PSF that comes from this object (fracflux¿0.5) to ensure that the detections are not subject to significant crowding, and thus have reliable astrometry and (less relevant for us) photometry. After initial experimenting we decided not to perform any filtering on “quality factor” (qf) field or flags derived either from unWISE coadds themselves (flags_unwise) or from crowdsource processing (flags_info) as they prevent most brighter objects from being detected, while not reducing significantly the amount of typical artefacts among the candidates. The positional accuracy of positions in epochal catalogues strongly depends on the source brightness. Our initial experiments demonstrated that crowdsource-measured uncertainties (dx and dy fields) adequately reflect it for objects fainter than 8th mag, while brighter ones show significantly larger scatter, most probably due to being saturated. Therefore, we assumed constant positional uncertainty of 1′′ for the latters. Moreover, we detected systematic bias between positions measured during forward and backward scans. To compensate for it, we added an extra positional error of 0.15′′ to all measurement uncertainties.
@@ 2.1 Motion detection algorithm
Name | Description | Importance |
pm | Estimated proper motion | 0.01 |
e_pm | Statistical error of estimated proper motion | 0.17 |
N | Number of epochs with detections | 0.10 |
corr_30 | Pearson correlation coefficient of the coordinates vs time for all points within 3′′ from expected position at different epochs | 0.23 |
corr | Pearson correlation coefficient of the coordinates vs time for all points within 3 brightness dependent coordinate uncertainty from expected position at different epochs | 0.16 |
magstd | Actual RMS of individual magnitude measurements in band | 0.14 |
magerr | Mean error of individual magnitude measurements in band | 0.06 |
mag | Mean magnitude in W2 band | 0.05 |
magchi2 | Reduced of individual magnitude measurements in band | 0.02 |
b | Galactic latitude | 0.01 |
NNNN1 | Number of detections with 1 or more catalogue points not belonging to the track within its 3 positional uncertainty | 0.02 |
NNNN3 | the same for 3 or more points | 0.01 |
NNNN2 | the same for 2 or more points | 0.01 |
NNNN4 | the same for 4 or more points | 0.001 |
NNNN5 | the same for 5 or more points | 0.001 |
Actual motion detection is performed using the algorithm similar to the one implemented in e.g. Fasbender & Nidever (2021) and based on finding clusters of points located along the lines in a three-dimensional space of sky coordinates and time. However, in contrast to optical data in Fasbender & Nidever (2021), it is enormously complicated by significant crowding and relatively low positional accuracy of the catalogue points, comparable to the actual proper motion of most objects over ten years time span.
@@
After applying various quality cuts as described above we analyzed the sets of all objects from epochal catalogues for every tile in order to find candidate moving objects. To do so, we converted their sky coordinates to standard coordinates relative to the tile center, and proceeded to detection of linear structures in resulting three-dimensional space (two angular coordinates plus time) using a simple “opportunistic” algorithm that consider all pairs of points from different epochs that satisfy the criteria of being separated by at least 3′′ (slightly more than 1 pixel in unWISE coadds, and slightly less than half of full-width at half maximum (FWHM) of its point-spread function (PSF)) but no more than 50′′ (for reducing the computational complexity). This criterion ensures significant motion of the candidates, effectively placing a lower limit on the detectable proper motion of about 0.3′′/year. In order to lower the contamination due to stationary objects, we excluded the pairs where either of the points has other detections within 1′′ from it in more than of epochs with the same brightness, within measurement uncertainties. Then, for every pair, we formed a candidate track, computed its expected positions for all the epochs available for that tile, and then selected the objects from corresponding epochs (this time without applying any additional criteria on top of initial quality cuts) that are closer than 3 positional uncertainty to these positions. We rejected the candidate tracks that do not have measurements in more than 60% of the epochs, corresponding to at least 10 epochs in typical 16-epochal tile without low-quality epochs. We also required the modulo of Pearson Correlation Coefficient of at least one angular coordinate with time to be at least 0.5. For the tracks passing these criteria, we computed the proper motion by performing robust linear regression of their standard coordinates versus time using RANSAC(Fischler & Bolles, 1981) algorithm as implemented in scikit-learn(Pedregosa et al., 2011) Python package in order to mitigate the influence of spurious associations with stationary objects, and then rejected the tracks with relative accuracy of proper motion estimation less than 0.33, thus ensuring that the candidate motion is significant. At this stage we also computed various statistical properties for the points belonging to the tracks after RANSAC filtering, as well as nearby catalogue points from different epochs, to serve later as features for artefact rejection.
@@
The candidates proposed by the algorithm are prone to both genuine image-level artefacts, especially due to long spikes around bright stars, and spurious tracks formed from individual detections of stationary objects located approximately along a line, especially for fainter candidates with larger positional uncertainties of individual detections. We could not find a way to reliably filter out either kind of artefacts using the information available in the catalogue files, and had to rely on the machine learning based filtering method that will be described in more details below, and finally on the visual vetting of the candidates with a simple Jupyter-based (Beg et al., 2021) dashboard based on the WISEView (Meisner et al., 2018; Caselden et al., 2018) cutout API developed in the frame of Backyard Worlds project (Kuchner et al., 2017). The dashboard displays, for every candidate, the cutouts from the first and last epochs at the object position, and the corresponding differences between these images and the median image of all epochs, and allows easy selection of real moving objects by the characteristic dipoles in the latter images, as shown in Figure 1. The dashboard displays up to several hundreds of candidates in a single page along with interactive controls for their flagging as either artefacts or actual moving objects, with the ability to also directly check the epochal animations using the original WISEView web tool111Available at http://byw.tools/wiseview. in the unclear cases.
The algorithm returned 1,271,921 initial candidates from all tiles of unTimely catalogue. Despite the visual vetting dashboard allowing a person to inspect thousands of the candidates in relatively small amount of time, this amount is still too much. Therefore we developed a simple iterative routine, akin to the one used by Kota et al. (2022), for extracting “true” moving objects from them based on Random Forest222We used Random Forest classifier implementation from scikit-learn (Pedregosa et al., 2011), with 100 estimators and automatically balanced class weights. (Breiman, 2001) binary classifier and a set of features for the candidates extracted by the algorithm, as listed in Table 1. We bootstrapped the classifier by selecting and visually checking the candidates with 60 largest (which all appeared to be true moving objects) and 60 smallest (which all were artefacts) values of the correlation of coordinates with time (corr parameter), and training the Random Forest on these objects. We then applied the classifier to all candidates, selected a hundred ones with largest predicted probability of being true, and visually inspected them, then using derived labels to improve the classifier and repeat the whole process. It’s progress is shown in Figure 2. We finished the process when the classifier stopped predicting new candidates with non-zero probabilities. At that point, we got 21,885 candidates visually vetted to be real moving objects, and 37,437 – as various artefacts. We characterized the performance of final classifier using Leave One Out cross-validation – by repeatedly excluding single points to serve as a test set, and training the model on the rest. After doing full cycle, we estimated mean precision of the classifier to be 99.6%, and mean recall – 96.5%.
Figure 3 shows the distribution of several features for both real moving objects and artefacts, both visually vetted and rejected by the classifier. The amount of the latters significantly grows towards lower number of epochs, larger uncertainties of estimated proper motions, larger photometric scatter, as well as towards the Galactic plane with largest density of stationary objects.
2.2 Performance verification
As a result of iterative machine learning based filtering of initial candidates described above, we visually confirmed 21,885 moving objects among 59,322 candidates we manually inspected in total, giving 37% overall efficiency of the algorithm as a whole. In order to control the sensitivity of the algorithm to different values of proper motion we performed a run of the candidate extraction routine with additional injection into every tile of 100 simulated objects with random positions, velocities and brightness, and positional scatter corresponding to actual uncertainties for this brightness, as described above. We processed these simulated objects together with actual data, so that they have the same density of background distractor points as real moving objects, and checked the fraction of these events successfully recovered by the algorithm. As expected (see Figure 4), the algorithm practically does not detect the objects moving slower than 0.3′′/year, but recovers all of the ones moving as fast as 10′′/year (corresponding to fastest known star), and also most of the ones up to 20′′/year, thus proving that our choice of the algorithm parameters is sufficient for detecting even fastest of moving objects at stellar distances.
As we simulated the data on catalogue level, we expect that this test does not include all selection effects affecting real data. Moreover, it does not include the machine learning classification step (as simulation does not realistically reproduce all the data properties needed for computing the features). Therefore, we also validated the algorithm by cross-matching moving objects we detected with the UltracoolSheet (Best et al., 2024, 2021, 2024) compilative database of ultracool dwarfs333UltracoolSheet is available online at http://bit.ly/UltracoolSheet. that contains the most comprehensive list of known nearby cool objects. We filtered it to only contain entries having WISE associations to ensure they are potentially visible, and checked the fraction that is actually detected by the algorithm and passed machine learning filtering, as described above. As shown in Figure 4, the efficiency estimated this way is consistent with simulations for brighter objects () and is slightly less for fainter ones due to unTimely epochal catalogues being not as deep as CatWISE2020 based on full unWISE coadds (Marocco et al., 2021).
To assess the accuracy of the proper motion determination by the algorithm we performed a cross-match of detected objects with Gaia DR3 catalogue (Gaia Collaboration et al., 2023) by propagating object positions to its J2016.0 epoch, using 2′′ matching radius, and only considering Gaia objects with significant proper motions to avoid confusions with faint stationary background stars. This way we found the associations for 20555 (94%) of all candidates. For 104 more candidates, mostly brighter ones, where Gaia DR3 did not provide an association we found matching counterparts in Gaia DR2 catalogue. Figure 5 shows that the agreement of proper motions is sufficiently good and unbiased for its whole range, except for the slowest objects with shortest arcs where our measurements tend to be overestimated so that they pass the 0.3′′/year algorithm threshold.